OpenStudy (anonymous):

If the range of the function f(x) = 4 − 4x is {-20, -16, -8, 0, 4}, what is its domain? A{0, 1, 2, 4, 6} B{2, 3, 5, 6, 7} C{1, 3, 5, 7, 9} D{0, 1, 3, 5, 6}

jimthompson5910 (jim_thompson5910):

Pick a value out of the range, say 4 plug 4 into f(x) to get f(x) = 4 − 4x 4 = 4 − 4x now solve for x to get x = ??

OpenStudy (anonymous):

I still don't get it...I plugged the 4 in for f(x) but got 21.

jimthompson5910 (jim_thompson5910):

solve 4 = 4 − 4x for x and tell me what you get

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

so if you were to plug x = 0 into f(x) = 4 − 4x, you'd get 4 when x = 0, f(x) = 4

jimthompson5910 (jim_thompson5910):

I'm going backwards because they give you the range (which is the set of y values or outputs)

jimthompson5910 (jim_thompson5910):

so you have to figure out which x values give you those specific y values

OpenStudy (anonymous):

I still honestly don't understand..

jimthompson5910 (jim_thompson5910):

The range is {-20, -16, -8, 0, 4}

jimthompson5910 (jim_thompson5910):

there are specific values of x that produce those y values

jimthompson5910 (jim_thompson5910):

when x = 0, y is y = 4 − 4x y = 4 − 4(0) y = 4 - 0 y = 4 So the input of x = 0 produces the output y = 4

jimthompson5910 (jim_thompson5910):

What is the input when the output is y = -8 ?